Lan Co. had 250,000 shares of common stock issued and outstanding at January 1, 20X6. During 20X6, Lan took the following actions:
March 15—Declared a 3-for-1 stock split, when the fair value of the stock was $80 per share.
December 15—Declared a $1.50 per share cash dividend.
In Lan's statement of stockholders' equity for 20X6, what amount should Lan report as dividends?
A. $1,125,000
B. $750,000
C. $375,000
D. $250,000
On May 18, 20X4, Son Corp.'s board of directors declared a 20% stock dividend.
The market price of Son's 5,000 outstanding shares of $4 par value common stock was $13 per share on that date. The stock dividend was distributed on July 21, 20X4, when the stock's market price was $15 per share.
What amount should Son credit to additional paid-in capital for this stock dividend?
A. $45,000
B. $4,500
C. $11,000
D. $9,000
The owners' equity section of a firm includes (1) $20,000 of 9%, $100 par cumulative preferred stock, and (2) $60,000 of $5 par common stock. There is additional paid-in capital on both issues. The preferred is fully participating and there are two years of dividends in arrears as of the beginning of the current year. If the firm pays $50,000 in dividends, what amount is allocated to common?
A. $50,000
B. $15,200
C. $34,800
D. $0
Answer #1

(1) Common Stock before stock split = 250000 shares

        Common stock after stock split = 250000 * 3 = 750000 shares

Amount of Dividend = 750000 * $1.5 = $1125000

Option A is correct

(2) Additional Paid in capital for stock dividend

Stock Dividend shares = 5000 * 20% = 1000 shares

Market Price on the date of declaration = $13

Additional Paid in capital = 13 – 4 = $9

$9 * 1000 shares = $9000

Option D is correct

(3) Amount allocated to common shareholders:-

#Annual Preferred= already give $20000#
#Annual Common = Already give $60000
Total 80000, Pref=1/4 Comm= 3/4
Pref, two years arrear
1. Arrears = P ($20000 * 9% * 2 years) 3600
2. Current = P ($20000 * 9%) 1800
3. Matching =C(60000* 9%) 5400
Until now, 50000-3600-1800-5400=39200
4. Parti P (1/4 * 39200) 9800
5. Parti Comm (3/4 *39200) 29400

Comm= 5400 + 29400 = 34800

Option C is correct